# Lense-Thirring Acoustic Black Holes : Shadows and Light

### v1

@@ Abstract
1 Département de Physique, Equipe des Sciences de la matière et du rayonnement, ESMAR
Faculté des Sciences, Université Mohammed V de Rabat, Rabat, Morocco
anas.elbalali@gmail.com
2 Centre for Mathematics and Theoretical Physics,
University of Hertfordshire, AL10 9AB Hatfield, UK
a.marrani@herts.ac.uk
@@ 1 Introduction
In Einstein’s theory of General Relativity, the presence of matter in the universe dictates how the spacetime shapes, as described by Einstein’s field equations [1]. Among the solutions to these equations, we find black holes. These objects are the result of the gravitational collapse of massive stars, at the end of their life cycle [2, 3]. Such a phenomenon leads to the formation of a singularity, surrounded by the event horizon, ‘the sphere of no return’. Black holes are therefore characterized by a very intense gravity that influences the surrounding spacetime, and which may thus lead to interesting optical phenomena. Indeed, it can be shown that black holes can act as a lens, causing the deflection of light (gravitational lensing) [4, 5, 6, 7, 8, 9, 10, 11]. From another perspective, the so-called black hole shadow is also a consequence of the immense gravitational pull of these objects [12, 13, 14, 15, 16, 17].
@@
Furthermore, black holes exhibit thermodynamic proprieties such as the entropy, the temperature, and most importantly the Hawking radiation, in which quantum effects are responsible for the black hole to emit radiation and slowly evaporate [18, 19, 20, 21, 22, 23, 24, 25]. Along this and other research venues, the study of black holes keep providing remarkable insights on the unification of quantum mechanics and gravity [26, 27, 28, 29, 30]. Black holes have also played an important role in the understanding of gravitational waves : the detection of GW 150914GW\,150914 published by LIGO and Virgo Collaboration has marked the first gravitational wave detection resulting from the merger of a binary black hole system [31]. It should also be recalled that imaging by the Event Horizon Telescope have provided an undeniable evidence of the existence of black holes themselves [32, 33, 34, 35, 36, 37, 38, 39]. Such discoveries and unprecedented advances in the observation of black holes have aroused the interest of physicists around the world leading to a relentless and ever growing research activity.
@@
Notwithstanding such ground-breaking results in the field of black hole physics, the experimental detection of the black hole features is still subject to limiting technical limitations. This can be regarded as one of the motivations which led to the birth of a new field of research in gravity, the so-called analogue gravity, aiming at mimicking gravitational pheonomena in the laboratory; needless to say, analogies have been constantly playing a key role in mathematics and physics since they have offered new perspectives and simplifications of complex ideas. In particular, it is here worth remarking that, through table top experiments, analogue gravity established key connections between rotating black holes and the Hawking radiation. This was firstly reported in the seminal paper by Unruh, in which hydrodynamical flows were used to mimic some proprieties of black holes [40]. In this framework, sound waves which are trapped inside a horizon led to the introduction of the name acoustic black hole or dumb hole. All in all, it has been argued that any fluid can ultimately give birth to an acoustic black hole, if it moves faster than the local sound velocity within a spherical surface [41, 42]. Black hole analogues have been investigated within a wide range of classical and quantum systems; just to cite a few, gravity waves [43], water [44], ‘slow’ light [45, 46, 47], optical fibers [48], and electromagnetic waveguides [49]. Other models exhibiting promising features include the superfluid helium II [50], atomic Bose-Einstein condensates [51], and one-dimensional Fermi-degenerate non-interacting gases [52]. For further detail and a quite complete list of references, we address the interested reader to the comprehensive review [53].
@@
Acoustic black holes have long been the subject of many studies. In a thermodynamical framework, Hawking-like radiation emitted by acoustic black holes has been successfully detected and studied in [54, 55, 56], and entropy of acoustic black holes has also been investigated in a number of studies [57, 58, 59, 60, 61]. More insights were gained by focussing on two-dimensional systems [62]. in which the investigation of the Bose-Einstein condensates provided an analogue of the Bekenstein-Hawking entropy [63]. While in [64] the entanglement entropy of acoustic black holes was investigated, the relation between the area of the acoustic black hole event horizon an the entropy has been the subject of [65]. Theoretically, the consideration of relativistic Gross-Pitaevski and Yang-Mills theories led to the embedding of the acoustic black hole into a curved background [66]; for instance, in [67] an acoustic black hole solution was related to a black $D3$-brane. Moreover, the investigation of such black hole analogues in the Abelian Higgs model led to the discovery of a relativistic acoustic black hole solution in Minkowski spacetime [68, 69].
@@
An interesting property of acoustic black holes goes under the name of acoustic shadow : it can be regarded as the acoustic analogue of the optical shadow of a black hole that has been captured by the EHT collaboration, and it has been studied in a variety of backgrounds and contexts; see e.g. [70, 71, 72]. The deflection angle has also been scrutinised in proximity of an acoustic black hole, in [73, 74]. At any rate, realistic models of acoustic black holes must necessarily include rotation : this is one of the motivations of [75], in which a slowly rotating solution has been constructed and investigated. In this framework, the interest of the analogue gravity approach towards the relativistic Gross-Pitaevskii theory as well as towards Yang-Mills (and thus, non-Abelian) gauge theories relies on a number of reasons; here, we will list three of them. Firstly, realistic models of astrophysical black holes include the embedding into the cosmological microwave background, and this makes the acoustic black holes natural candidates to be investigated in this venue [66]. Secondly, the relativistic and transonic accretion disks, surrounding supermassive black holes at the center of galaxies, serve as unique examples of analogue gravity models present in nature [76, 77, 78, 79]. Thirdly, one shoud recall that black holes may also be surrounded by some quantum superfluids whose condensation can generally give rise to analogue gravitational phenomena; for instance, such fluids have been proposed to determine also dark matter phenomenology [80].
@@ 3 Review of Lense-Thirring acoustic black hole solution
In this part of the paper, we review the procedure to derive a slowly rotating acoustic black hole solution. In order to establish this solution, we follow the approach outlined in paper [75]. Indeed, within the framework of analogue gravity, the Lagrangian of relativistic Gross-Pitaevskii (GP) theory of a complex scalar field $\varphi$ is written as
$\mathcal{L}=g^{\mu\nu}\partial_{\mu}\varphi^{\star}\partial_{\nu}\varphi-m^{2}\varphi^{\star}\varphi+\frac{b}{2}\left(\varphi^{\star}\varphi\right)^{2},$ (3.1)
@@
$S=\int d^{4}x\sqrt{-g}\,\left(\big|\partial_{\mu}\varphi\big|^{2}-m^{2}\big|\varphi\big|^{2}+\frac{b}{2}\big|\varphi\big|^{4}\right),$ (3.2)
@@
$\square\varphi+m^{2}\varphi-b\big|\varphi\big|^{2}\varphi=0.$ (3.3)
@@
$ds^{2}=-g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}+2g_{t\phi}dtd\phi.$ (3.4)
It is worth recalling that such a metric form has been chosen to align with a slowly rotating black hole, similar to the Lense-Thirring Black Hole (LTBH), which will be introduced later [82]. Indeed, the slowly rotating approximation assumes that higher-order terms of the rotation parameter $a$, which is embedded in the component $g_{t\phi}$, are negligible. The scalar field $\varphi$ can be expressed as a function of the fluid density $\rho=\rho_{0}+\epsilon\rho_{1}$ and the phase $\vartheta=\vartheta_{0}+\epsilon\vartheta_{1}$ as follows
$\varphi=\sqrt{\rho\left(\vec{x},t\right)}\exp\left(i\vartheta\left(\vec{x},t\right)\right).$ (3.5)
In the hydrodynamic or Thomas-Fermi limit the quantum pressure $P_{\mathcal{Q}}=\frac{\square\sqrt{\rho}}{\sqrt{\rho}}$ can be neglected in front of the quantities $m^{2},b\rho_{0},v^{\mu}v_{\mu}$. Thus, replacing $\varphi$ in equation (3.3), one finds the background fluid density leading order equation which reads
$b\rho_{0}\equiv 2c_{s}^{2}=m^{2}-g^{\mu\nu}v_{\mu}v_{\nu},$ (3.6)
@@
$\nabla_{\mu}\left(\rho v^{\mu}\right)=\nabla_{\mu}\left(\rho_{0}v_{0}^{\mu}\right)+\epsilon\nabla_{\mu}\left(\rho_{0}\partial^{\mu}\vartheta_{1}+\rho_{1}v_{0}^{\mu}\right)=0,$ (3.7)
@@
$\nabla_{\mu}\left(\rho_{0}\partial^{\mu}\vartheta_{1}+\rho_{1}v_{0}^{\mu}\right)=0.$ (3.8)
At subleading order in $\epsilon$ [81], the decomposition of the scalar field $\varphi$ when replaced in equation of motion of the scalar field (3.3) leads to
$\rho_{1}=-\frac{2}{b}v_{0}^{\mu}\partial_{\mu}\vartheta_{1}.$ (3.9)
Replacing in equation (3.8), one gets
$\nabla_{\mu}\left(\rho_{0}g^{\mu\nu}-\frac{2}{b}v_{0}^{\mu}v_{0}^{\nu}\right)\partial_{\nu}\vartheta_{1}=0.$ (3.10)
@@
$\mathcal{G}^{\mu\nu}\propto\rho_{0}g^{\mu\nu}-\frac{2}{b}v_{0}^{\mu}v_{0}^{\nu}.$ (3.11)
@@
$\frac{1}{\sqrt{-\mathcal{G}}}\partial_{\mu}\left(\sqrt{-\mathcal{G}}\,\mathcal{G}^{\mu\nu}\partial_{\nu}\vartheta_{1}\right)=0.$ (3.12)
@@
$dt$ $\rightarrow dt+\frac{v_{r}v_{t}}{g_{tt}\left(c_{s}^{2}-v_{i}v^{i}\right)}dr,$ (3.13)
$d\phi$ $\rightarrow d\phi-\frac{g_{t\phi}\left(c_{s}^{2}-v_{t}v^{t}\right)v_{t}v_{r}}{g_{\phi\phi}g_{tt}\left(c_{s}^{2}-v_{\mu}v^{\mu}\right)\left(c_{s}^{2}-v_{r}v^{r}\right)}dr,$ (3.14)
@@
$ds^{2}=c_{s}\sqrt{c_{s}^{2}-v_{\mu}v^{\mu}}\left(\frac{c_{s}^{2}-v_{r}v^{r}}{c_{s}^{2}-v_{\mu}v^{\mu}}g_{tt}dt^{2}+\frac{c_{s}^{2}}{c_{s}^{2}-v_{r}v^{r}}g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}+2\frac{c_{s}^{2}-v_{t}v^{t}}{c_{s}^{2}-v_{\mu}v^{\mu}}g_{t\phi}dtd\phi\right).$ (3.15)
@@
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}-2\left(1-f(r)\right)a\sin^{2}\theta dtd\phi,$ (3.16)
@@
$f(r)=1-\frac{2M}{r},$ (3.17)
and $a$ is the rotation parameter [82]. Regarding the four-velocity, one can use the equation (3.6) and rescale in terms of natural units of $2c_{s}^{2}$ as follows
$\tilde{v}_{\mu}\rightarrow\frac{v_{\mu}}{\sqrt{2c_{s}^{2}}},$ (3.18)
@@
$\tilde{v}_{\mu}\tilde{v}^{\mu}=\frac{m^{2}}{2c_{s}^{2}}-1.$ (3.19)
@@
$v_{r}\sim\sqrt{\left(1-f(r)\right)\xi},$ (3.21)
@@
$v_{\mu}v^{\mu}=g^{tt}v_{t}^{2}+g^{rr}v_{r}^{2}=-\frac{v_{t}^{2}}{f(r)}+f(r)v_{r}^{2}.$ (3.22)
With $v_{\mu}v^{\mu}=-1$, equation (3.22) gives
$v_{t}=\sqrt{f(r)+\left(1-f(r)\right)f(r)^{2}\xi}.$ (3.23)
Finally, the metric of equation (3.15) can be rewritten as
$ds^{2}=-F(r)dt^{2}+\frac{dr^{2}}{F(r)}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}-2\left(1-F(r)\right)a\sin^{2}dtd\phi,$ (3.24)
@@
$F(r)=f(r)\left[1-f(r)\left(1-f(r)\right)\xi\right].$ (3.25)
For simplicity reasons, the speed of sound has been fixed, i.e $c_{s}^{2}=1/\sqrt{3}$. It is worth noticing that, the escape velocity $v_{r}$ in equation (3.21) implies that the parameter $\xi$ must be strictly positive. Besides, when the acoustic parameter $\xi\rightarrow 0$, we recover the LTBH while when $\xi\rightarrow\infty$, i.e $v_{r}\rightarrow\infty$, the entire spacetime is covered by the LTABH.
The spacetime geometry of a LTABH is described by equations (3.24), (3.25), and f(r)=12M/rf(r)=1-2M/r. Such scenario offers valuable insight which deserves further investigation. In what follows, we analyse the geometry of the this solution, discuss the roots of the acoustic metric function $F(r)$ in terms of the acoustic parameter $\xi$ and derive the shadow equation. We also investigate the critical radii involved in the acoustic case and shed light on the frame dragging effect and the deflection angle of the present solution.

### v2

# Faculté des Sciences, Université Mohammed V de Rabat, Rabat, Morocco

@@ 1 Introduction
In Einstein’s theory of General Relativity, the presence of matter in the universe dictates how the spacetime shapes, as described by Einstein’s field equations Ip11 . Among the solutions to these equations, we find black holes. These objects are the result of the gravitational collapse of massive stars, at the end of their life cycle Ip12 ; Ip13 . Such a phenomenon leads to the formation of a singularity, surrounded by the event horizon, ‘the sphere of no return’. Black holes are therefore characterized by a very intense gravity that influences the surrounding spacetime, and which may thus lead to interesting optical phenomena. Indeed, it can be shown that black holes can act as a lens, causing the deflection of light (gravitational lensing) Ip14 ; Ip15 ; Ip16 ; Ip17 ; Ip18 ; Ip19 ; Ip110 ; Ip111 . From another perspective, the so-called black hole shadow is also a consequence of the immense gravitational pull of these objects Ip112 ; Ip113 ; Ip114 ; Ip115 ; Ip116 ; Ip117 .
@@
Furthermore, black holes exhibit thermodynamic proprieties such as the entropy, the temperature, and most importantly the Hawking radiation, in which quantum effects are responsible for the black hole to emit radiation and slowly evaporate Ip118 ; Ip119 ; Ip120 ; Ip121 ; Ip122 ; Ip123 ; Ip124 ; Ip125 . Along this and other research venues, the study of black holes keep providing remarkable insights on the unification of quantum mechanics and gravity Ip126 ; Ip127 ; Ip128 ; Ip129 ; Ip130 . Black holes have also played an important role in the understanding of gravitational waves : the detection of GW 150914GW\,150914 published by LIGO and Virgo Collaboration has marked the first gravitational wave detection resulting from the merger of a binary black hole system Ip131 . It should also be recalled that imaging by the Event Horizon Telescope have provided an undeniable evidence of the existence of black holes themselves Ip132 ; Ip133 ; Ip134 ; Ip135 ; Ip136 ; Ip137 ; Ip138 ; Ip139 . Such discoveries and unprecedented advances in the observation of black holes have aroused the interest of physicists around the world leading to a relentless and ever growing research activity.
@@
Notwithstanding such ground-breaking results in the field of black hole physics, the experimental detection of the black hole features is still subject to limiting technical limitations. This can be regarded as one of the motivations which led to the birth of a new field of research in gravity, the so-called analogue gravity, aiming at mimicking gravitational pheonomena in the laboratory; needless to say, analogies have been constantly playing a key role in mathematics and physics since they have offered new perspectives and simplifications of complex ideas. In particular, it is here worth remarking that, through table top experiments, analogue gravity established key connections between rotating black holes and the Hawking radiation. This was firstly reported in the seminal paper by Unruh Ip21 , in which hydrodynamical flows were used to mimic some proprieties of black holes. In this framework, sound waves which are trapped inside a horizon led to the introduction of the name acoustic black hole or dumb hole. All in all, it has been argued that any fluid can ultimately give birth to an acoustic black hole, if it moves faster than the local sound velocity within a spherical surface Ip22 ; Ip23 ; If1 ; If2 . Black hole analogues have been investigated within a wide range of classical and quantum systems; just to cite a few, gravity waves Ip24 , water Ip25 , ‘slow’ light Ip26 ; Ip27 ; Ip28 , optical fibers Ip29 , and electromagnetic waveguides Ip210 . Other models exhibiting promising features include the superfluid helium II Ip211 , atomic Bose-Einstein condensates Ip212 , and one-dimensional Fermi-degenerate non-interacting gases Ip213 . For further detail and a quite complete list of references, we address the interested reader to the comprehensive review Ip214 .
@@
Acoustic black holes have long been the subject of many studies. In a thermodynamical framework, Hawking-like radiation emitted by acoustic black holes has been successfully detected and studied in Ip215 ; Ip216 ; Ip217 , and entropy of acoustic black holes has also been investigated in a number of studies Ip2171 ; Ip2172 ; Ip2173 ; Ip2174 ; Ip2175 . More insights were gained by focussing on two-dimensional systems Ip2176 ; Ip21761 , in which the investigation of the Bose-Einstein condensates provided an analogue of the Bekenstein-Hawking entropy Ip2177 . While in Ip2178 the entanglement entropy of acoustic black holes was investigated, the relation between the area of the acoustic black hole event horizon an the entropy has been the subject of Ip2179 . Theoretically, the consideration of relativistic Gross-Pitaevski and Yang-Mills theories led to the embedding of the acoustic black hole into a curved background Ip218 ; for instance, in Ip219 an acoustic black hole solution was related to a black $D3$-brane. Moreover, the investigation of such black hole analogues in the Abelian Higgs model led to the discovery of a relativistic acoustic black hole solution in Minkowski spacetime Ip22 ; Ip221 .
@@
An interesting property of acoustic black holes goes under the name of acoustic shadow : it can be regarded as the acoustic analogue of the optical shadow of a black hole that has been captured by the EHT collaboration, and it has been studied in a variety of backgrounds and contexts; see e.g. Ip32 ; Ip33 . The deflection angle has also been scrutinised in proximity of an acoustic black hole, in Ip34 ; Ip35 . At any rate, realistic models of acoustic black holes must necessarily include rotation : this is one of the motivations of main , in which a slowly rotating solution has been constructed and investigated. In this framework, the interest of the analogue gravity approach towards the relativistic Gross-Pitaevskii theory as well as towards Yang-Mills (and thus, non-Abelian) gauge theories relies on a number of reasons; here, we will list three of them. Firstly, realistic models of astrophysical black holes include the embedding into the cosmological microwave background, and this makes the acoustic black holes natural candidates to be investigated in this venue Ip218 . Secondly, the relativistic and transonic accretion disks, surrounding supermassive black holes at the center of galaxies, serve as unique examples of analogue gravity models present in nature Ip331 ; Ip332 ; Ip333 ; Ip334 . Thirdly, one shoud recall that black holes may also be surrounded by some quantum superfluids whose condensation can generally give rise to analogue gravitational phenomena; for instance, such fluids have been proposed to determine also dark matter phenomenology Ip335 .
@@ 3 Review of Lense-Thirring acoustic black hole solution
In this part of the paper, we review the procedure to derive a slowly rotating acoustic black hole solution. In order to establish this solution, we follow the approach outlined in paper main . Indeed, within the framework of analogue gravity, the Lagrangian of relativistic Gross-Pitaevskii (GP) theory of a complex scalar field $\varphi$ is written as
$\mathcal{L}=g^{\mu\nu}\partial_{\mu}\varphi^{\star}\partial_{\nu}\varphi-m^{2}\varphi^{\star}\varphi+\frac{b}{2}\left(\varphi^{\star}\varphi\right)^{2},$ (1)
@@
$S=\int d^{4}x\sqrt{-g}\,\left(\big|\partial_{\mu}\varphi\big|^{2}-m^{2}\big|\varphi\big|^{2}+\frac{b}{2}\big|\varphi\big|^{4}\right),$ (2)
@@
$\square\varphi+m^{2}\varphi-b\big|\varphi\big|^{2}\varphi=0.$ (3)
@@
$ds^{2}=-g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}+2g_{t\phi}dtd\phi.$ (4)
It is worth recalling that such a metric form has been chosen to align with a slowly rotating black hole, similar to the Lense-Thirring Black Hole (LTBH), which will be introduced later main3 . Indeed, the slowly rotating approximation assumes that higher-order terms of the rotation parameter $a$, which is embedded in the component $g_{t\phi}$, are negligible. The scalar field $\varphi$ can be expressed as a function of the fluid density $\rho=\rho_{0}+\epsilon\rho_{1}$ and the phase $\vartheta=\vartheta_{0}+\epsilon\vartheta_{1}$ as follows
$\varphi=\sqrt{\rho\left(\vec{x},t\right)}\exp\left(i\vartheta\left(\vec{x},t\right)\right).$ (5)
In the hydrodynamic or Thomas-Fermi limit the quantum pressure $P_{\mathcal{Q}}=\frac{\square\sqrt{\rho}}{\sqrt{\rho}}$ can be neglected in front of the quantities $m^{2},b\rho_{0},v^{\mu}v_{\mu}$. Thus, replacing $\varphi$ in equation (3), one finds the background fluid density leading order equation which reads
$b\rho_{0}\equiv 2c_{s}^{2}=m^{2}-g^{\mu\nu}v_{\mu}v_{\nu},$ (6)
@@
$\nabla_{\mu}\left(\rho v^{\mu}\right)=\nabla_{\mu}\left(\rho_{0}v_{0}^{\mu}\right)+\epsilon\nabla_{\mu}\left(\rho_{0}\partial^{\mu}\vartheta_{1}+\rho_{1}v_{0}^{\mu}\right)=0,$ (7)
@@
$\nabla_{\mu}\left(\rho_{0}\partial^{\mu}\vartheta_{1}+\rho_{1}v_{0}^{\mu}\right)=0.$ (8)
At subleading order in $\epsilon$ main2 , the decomposition of the scalar field $\varphi$ when replaced in equation of motion of the scalar field (3) leads to
$\rho_{1}=-\frac{2}{b}v_{0}^{\mu}\partial_{\mu}\vartheta_{1}.$ (9)
Replacing in equation (8), one gets
$\nabla_{\mu}\left(\rho_{0}g^{\mu\nu}-\frac{2}{b}v_{0}^{\mu}v_{0}^{\nu}\right)\partial_{\nu}\vartheta_{1}=0.$ (10)
@@
$\mathcal{G}^{\mu\nu}\propto\rho_{0}g^{\mu\nu}-\frac{2}{b}v_{0}^{\mu}v_{0}^{\nu}.$ (11)
@@
$\frac{1}{\sqrt{-\mathcal{G}}}\partial_{\mu}\left(\sqrt{-\mathcal{G}}\,\mathcal{G}^{\mu\nu}\partial_{\nu}\vartheta_{1}\right)=0.$ (12)
@@
$dt$ $\rightarrow dt+\frac{v_{r}v_{t}}{g_{tt}\left(c_{s}^{2}-v_{i}v^{i}\right)}dr,$ (13)
$d\phi$ $\rightarrow d\phi-\frac{g_{t\phi}\left(c_{s}^{2}-v_{t}v^{t}\right)v_{t}v_{r}}{g_{\phi\phi}g_{tt}\left(c_{s}^{2}-v_{\mu}v^{\mu}\right)\left(c_{s}^{2}-v_{r}v^{r}\right)}dr,$ (14)
@@
$ds^{2}=c_{s}\sqrt{c_{s}^{2}-v_{\mu}v^{\mu}}\left(\frac{c_{s}^{2}-v_{r}v^{r}}{c_{s}^{2}-v_{\mu}v^{\mu}}g_{tt}dt^{2}+\frac{c_{s}^{2}}{c_{s}^{2}-v_{r}v^{r}}g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}+2\frac{c_{s}^{2}-v_{t}v^{t}}{c_{s}^{2}-v_{\mu}v^{\mu}}g_{t\phi}dtd\phi\right).$ (15)
@@
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}-2\left(1-f(r)\right)a\sin^{2}\theta dtd\phi,$ (16)
@@
$f(r)=1-\frac{2M}{r},$ (17)
and $a$ is the rotation parameter main3 . Regarding the four-velocity, one can use the equation (6) and rescale in terms of natural units of $2c_{s}^{2}$ as follows
$\tilde{v}_{\mu}\rightarrow\frac{v_{\mu}}{\sqrt{2c_{s}^{2}}},$ (18)
@@
$\tilde{v}_{\mu}\tilde{v}^{\mu}=\frac{m^{2}}{2c_{s}^{2}}-1.$ (19)
@@
$v_{r}\sim\sqrt{\left(1-f(r)\right)\xi},$ (21)
@@
$v_{\mu}v^{\mu}=g^{tt}v_{t}^{2}+g^{rr}v_{r}^{2}=-\frac{v_{t}^{2}}{f(r)}+f(r)v_{r}^{2}.$ (22)
With $v_{\mu}v^{\mu}=-1$, equation (22) gives
$v_{t}=\sqrt{f(r)+\left(1-f(r)\right)f(r)^{2}\xi}.$ (23)
Finally, the metric of equation (15) can be rewritten as
$ds^{2}=-F(r)dt^{2}+\frac{dr^{2}}{F(r)}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}-2\left(1-F(r)\right)a\sin^{2}dtd\phi,$ (24)
@@
$F(r)=f(r)\left[1-f(r)\left(1-f(r)\right)\xi\right].$ (25)
For simplicity reasons, the speed of sound has been fixed, i.e $c_{s}^{2}=1/\sqrt{3}$. It is worth noticing that, the escape velocity $v_{r}$ in equation (21) implies that the parameter $\xi$ must be strictly positive. Besides, when the acoustic parameter $\xi\rightarrow 0$, we recover the LTBH while when $\xi\rightarrow\infty$, i.e $v_{r}\rightarrow\infty$, the entire spacetime is covered by the LTABH.
The spacetime geometry of a LTABH is described by equations (24), (25), and f(r)=12M/rf(r)=1-2M/r. Such scenario offers valuable insight which deserves further investigation. In what follows, we analyse the geometry of the this solution, discuss the roots of the acoustic metric function $F(r)$ in terms of the acoustic parameter $\xi$ and derive the shadow equation. We also investigate the critical radii involved in the acoustic case and shed light on the frame dragging effect and the deflection angle of the present solution.